a 250.0ml solution of 0.200mol L-1 aluminum nitrate was mixed together with 200.0ml of 0.200mol L-1 barium nitrate. What is the concentration of nitrate ions in the resultant solution?
1
Expert's answer
2020-06-08T15:44:21-0400

Moles of "NO^-_3" in aluminum nitrate"=3" times the moles of aluminum nitrate.(Chemical formula"\equiv Al(NO_3)_3" )

Moles of "NO_3^-" in aluminium nitrate solution"=3\times \frac{250}{1000}\times 0.200=0.15"

Moles of "NO_3^-" in barium nitrate"=2" times the moles of barium nitrate (Chemical formula"=Ba(NO_3)_2" )

Moles of "NO_3^-" in barium nitrate solution"=2\times \frac{200}{1000}\times 0.200= 0.08"

Total moles"=0.15+0.08=0.23"

Total volume"=200+250=450 \ ml=0.45 \ l"

Concentration in resultant solution"=\frac{0.23}{0.45}=0.51 \ mol/l"